# Finsler Geometry in Anisotropic Superconductivity: A Ginzburg–Landau Approach

### v1

@@ Abstract
Keywords: Finsler geometry; Ginzburg–Landau functional; anisotropic superconductivity; variational methods; $\Gamma$–convergence; vortex filaments.
[2020] Mathematics Subject Classification:53B40, 35J60, 35J20, 49J45, 58E50, 35Q56.
@@ 2 Preliminaries on Finsler Geometry
Proof.
Set $\xi_{t}:=d(u+t\varphi)=du+t\,d\varphi$. Using Proposition 2.1,
@@
Now apply Lemma 2.4 with $X=\nabla_{F}u$ to obtain
@@
which is equivalent to any other choice by Lemma 2.8.
Proposition 2.10.
@@
Proof.
By Lemma 2.8 and (7), the $H^{1}_{F}$–norm is equivalent to the standard $H^{1}(M,\gamma)$ norm with respect to the smooth positive density $\sigma\,dx$:
@@
For 11–forms we will use a fixed smooth Riemannian co-metric $\gamma^{*}$ uniformly equivalent to $F^{*}$ (Lemma 2.8) to formulate Hodge operators for the Maxwell term. Write $\sharp_{\gamma}:T^{*}M\to TM$ for the $\gamma$–musical isomorphism and define
@@
which follows by Lemma 2.4 and the definition of $\mathrm{div}_{\mu_{F}}$. In particular, $d^{\dagger}_{\gamma,\mu_{F}}$ is the $L^{2}(d\mu_{F})$–adjoint of $d$ acting on 0–forms.
@@ 3 The Finslerian Ginzburg–Landau Functional
Let $(M,F)$ be a compact smooth Finsler manifold endowed with a fixed smooth Finsler measure $d\mu_{F}$ (either Busemann–Hausdorff or Holmes–Thompson) and a smooth Riemannian co–metric $\gamma^{*}$ uniformly equivalent to $F^{*}$ (cf. Lemma 2.8). For a complex scalar field $\psi:M\to\mathbb{C}$ and a real $1$-form $A\in\Omega^{1}(M)$ we set
@@
Proof.
By Lemma 2.8 and compactness of $M$, $F^{*}(x,\cdot)$ is uniformly equivalent to $|\cdot|_{\gamma^{*}}$, hence $\|D_{A}\psi\|_{F^{*}}\in L^{2}(M,d\mu_{F})$ when $\psi,A\in H^{1}_{F}$. The Maxwell term is in $L^{1}$ because $dA\in L^{2}$ and $d\mu_{F}$ is smooth. The potential term is in $L^{1}$ since $H^{1}_{F}\hookrightarrow L^{4}$ on compact $M$ (Sobolev and norm equivalence). $C^{1}$-regularity follows from the chain rule and smoothness/convexity of $F^{*}(x,\cdot)^{2}$, plus bilinearity of $(\psi,A)\mapsto D_{A}\psi$. ∎
We compute the Gâteaux derivative of $G_{F}$ at $(\psi,A)$ in the directions $(\varphi,B)\in H^{1}_{F}(M;\mathbb{C})\times H^{1}_{F}(M;T^{*}M)$. By Definition 2.2, Theorem 2.7 (with complex realification), and the adjoint operator $d^{\dagger}_{\gamma,\mu_{F}}$, it follows that:
Proposition 3.4 (First variation).
@@
and apply Theorem 2.7 componentwise, noting that
@@
Proof.
By Lemma 2.8 and Proposition 2.10, $\|dA\|_{L^{2}}$ controls $\|A\|_{H^{1}}$ on $\mathcal{C}$ (standard elliptic estimate for the operator $d^{\dagger}d$ with Coulomb constraint). The kinetic term controls $\|\psi\|_{H^{1}_{F}}$ up to a constant via the diamagnetic inequality in the next lemma (and the potential term bounds $\|\psi\|_{L^{4}}$). Collect the bounds and absorb constants. ∎
Lemma 3.7.
@@ 4 Existence of Minimizers
be the natural energy space. Because $G_{F}$ is gauge invariant (Proposition 3.2), we restrict to a fixed Coulomb slice
@@
Proof.
Combine Lemma 3.6 (coercivity on Coulomb slice), Lemma 2.8 (uniform equivalence of $F^{*}$ and $\gamma^{*}$), and the Poincaré inequality (14). All constants depend only on $M,F,\gamma^{*},d\mu_{F}$ and $\lambda$. ∎
Corollary 4.3 (Bounded minimizing sequences).
@@
By the closedness of the Coulomb condition under weak convergence, $d^{\dagger}_{\gamma,\mu_{F}}A=0$, hence $(\psi,A)\in\mathcal{A}_{C}$. Applying Lemma 3.5 (sequential weak lower semi continuity),
@@
Theorem 4.4 (Existence of minimizers).
@@
Proof.
Existence follows directly from the compactness and l.s.c. arguments above. To verify the Euler–Lagrange equations, note that $G_{F}$ is Fréchet differentiable on $\mathcal{H}_{F}$ (Proposition 3.4), hence the first variation vanishes in all admissible directions in $\mathcal{A}_{C}$, yielding (17). ∎
@@
holds, where $\|D_{A_{\varepsilon}}\psi_{\varepsilon}\|_{L^{2}_{F}}^{2}:=\int_{M}\|D_{A_{\varepsilon}}\psi_{\varepsilon}\|_{F^{*}}^{2}\,d\mu_{F}$. Equation (19) is invariant under all gauge transformations due to Proposition 3.2.
When $F$ is Riemannian, i.e. $F(x,y)=\sqrt{g_{x}(y,y)}$, all definitions reduce to the classical ones for the magnetic Ginzburg–Landau model. The entire proof above specializes to the standard results of Bethuel–Brezis–Hélein [2] and Sandier–Serfaty [3]. Hence Theorem 4.4 can be viewed as their exact Finslerian extension.
@@ 5 Asymptotic Analysis and $\Gamma$–Convergence
Let us first recall that by the diamagnetic inequality (Lemma 3.7),
@@
Theorem 5.1 ($\Gamma$–convergence of $G_{F}$).
@@ 6 Numerical example
Discussion. The above analytic computations provide an explicit check that the Finslerian GL functional introduced in the paper recovers the expected anisotropic scaling in energy for simple topological configurations on the torus. The example is consistent with the general existence theory (Theorem 4.4) and with the $\Gamma$–-convergence statement (Theorem 5.1): energy concentration for sequences with growing winding or for configurations forcing zeros of $\psi$ would, after the standard $|\log\varepsilon|$–-rescaling, lead to limiting energies proportional to the Finsler length of the corresponding vortex currents; in this simple constant modulus family the energy is entirely carried by the phase gradient and computed above in closed form.
@@ 7 Concluding
In this work we have extended the classical Ginzburg–Landau theory to the setting of general Finsler manifolds. Starting from the analytic preliminaries in Section 2, we defined the anisotropic functional $G_{F}[\psi,A]$ in (16), proved its gauge invariance and well-posedness, and established the existence of minimizers (Theorem 4.4) by direct variational methods. Finally, in Section 5, we derived the full $\Gamma$–limit of the Finslerian energies as $\varepsilon\to 0$, showing that the limiting functional is the Finsler length of the vortex current.
@@
  • The Finsler diamagnetic inequality (Lemma 3.7) provides a geometric generalization of Kato’s inequality, valid for arbitrary convex co-metrics.
  • The variational proof of existence (Theorem 4.4) uses only the intrinsic Finsler Sobolev structure, avoiding Euclidean embeddings or local coordinates.
  • The $\Gamma$–limit (Theorem 5.1) identifies the limiting vortex energy with the anisotropic Finsler length $\int_{\Sigma}F(x,\nu_{\Sigma})\,d\mathcal{H}^{n-2}$, extending the isotropic theory of Bethuel–Brezis–Hélein and Sandier–Serfaty to arbitrary anisotropies.

### v2

# Open Communications in Nonlinear Mathematical Physics ]ocnmp[ Vol.5 (2025) pp 1References Article

@@ 2 Preliminaries on Finsler Geometry
Set $\xi_{t}:=d(u+t\varphi)=du+t\,d\varphi$. Using Proposition 1,
@@
Now apply Lemma 1 with $X=\nabla_{F}u$ to obtain
@@
which is equivalent to any other choice by Lemma 2.
Proposition 2.
@@
By Lemma 2 and (7), the $H^{1}_{F}$–norm is equivalent to the standard $H^{1}(M,\gamma)$ norm with respect to the smooth positive density $\sigma\,dx$:
@@
For 11–forms we will use a fixed smooth Riemannian co-metric $\gamma^{*}$ uniformly equivalent to $F^{*}$ (Lemma 2) to formulate Hodge operators for the Maxwell term. Write $\sharp_{\gamma}:T^{*}M\to TM$ for the $\gamma$–musical isomorphism and define
@@
which follows by Lemma 1 and the definition of $\mathrm{div}_{\mu_{F}}$. In particular, $d^{\dagger}_{\gamma,\mu_{F}}$ is the $L^{2}(d\mu_{F})$–adjoint of $d$ acting on 0–forms.
@@ 3 The Finslerian Ginzburg–Landau Functional
Let $(M,F)$ be a compact smooth Finsler manifold endowed with a fixed smooth Finsler measure $d\mu_{F}$ (either Busemann–Hausdorff or Holmes–Thompson) and a smooth Riemannian co–metric $\gamma^{*}$ uniformly equivalent to $F^{*}$ (cf. Lemma 2). For a complex scalar field $\psi:M\to\mathbb{C}$ and a real $1$-form $A\in\Omega^{1}(M)$ we set
@@
By Lemma 2 and compactness of $M$, $F^{*}(x,\cdot)$ is uniformly equivalent to $|\cdot|_{\gamma^{*}}$, hence $\|D_{A}\psi\|_{F^{*}}\in L^{2}(M,d\mu_{F})$ when $\psi,A\in H^{1}_{F}$. The Maxwell term is in $L^{1}$ because $dA\in L^{2}$ and $d\mu_{F}$ is smooth. The potential term is in $L^{1}$ since $H^{1}_{F}\hookrightarrow L^{4}$ on compact $M$ (Sobolev and norm equivalence). $C^{1}$-regularity follows from the chain rule and smoothness/convexity of $F^{*}(x,\cdot)^{2}$, plus bilinearity of $(\psi,A)\mapsto D_{A}\psi$. ∎
We compute the Gâteaux derivative of $G_{F}$ at $(\psi,A)$ in the directions $(\varphi,B)\in H^{1}_{F}(M;\mathbb{C})\times H^{1}_{F}(M;T^{*}M)$. By Definition Definition, Theorem 1 (with complex realification), and the adjoint operator $d^{\dagger}_{\gamma,\mu_{F}}$, it follows that:
Proposition 5 (First variation).
@@
and apply Theorem 1 componentwise, noting that
@@
By Lemma 2 and Proposition 2, $\|dA\|_{L^{2}}$ controls $\|A\|_{H^{1}}$ on $\mathcal{C}$ (standard elliptic estimate for the operator $d^{\dagger}d$ with Coulomb constraint). The kinetic term controls $\|\psi\|_{H^{1}_{F}}$ up to a constant via the diamagnetic inequality in the next lemma (and the potential term bounds $\|\psi\|_{L^{4}}$). Collect the bounds and absorb constants. ∎
Lemma 5.
@@ 4 Existence of Minimizers
be the natural energy space. Because $G_{F}$ is gauge invariant (Proposition 3), we restrict to a fixed Coulomb slice
@@
Combine Lemma 4 (coercivity on Coulomb slice), Lemma 2 (uniform equivalence of $F^{*}$ and $\gamma^{*}$), and the Poincaré inequality (14). All constants depend only on $M,F,\gamma^{*},d\mu_{F}$ and $\lambda$. ∎
Corollary 1 (Bounded minimizing sequences).
@@
By the closedness of the Coulomb condition under weak convergence, $d^{\dagger}_{\gamma,\mu_{F}}A=0$, hence $(\psi,A)\in\mathcal{A}_{C}$. Applying Lemma 3 (sequential weak lower semi continuity),
@@
Theorem 2 (Existence of minimizers).
@@
Existence follows directly from the compactness and l.s.c. arguments above. To verify the Euler–Lagrange equations, note that $G_{F}$ is Fréchet differentiable on $\mathcal{H}_{F}$ (Proposition 5), hence the first variation vanishes in all admissible directions in $\mathcal{A}_{C}$, yielding (17). ∎
@@
holds, where $\|D_{A_{\varepsilon}}\psi_{\varepsilon}\|_{L^{2}_{F}}^{2}:=\int_{M}\|D_{A_{\varepsilon}}\psi_{\varepsilon}\|_{F^{*}}^{2}\,d\mu_{F}$. Equation (19) is invariant under all gauge transformations due to Proposition 3.
When $F$ is Riemannian, i.e. $F(x,y)=\sqrt{g_{x}(y,y)}$, all definitions reduce to the classical ones for the magnetic Ginzburg–Landau model. The entire proof above specializes to the standard results of Bethuel–Brezis–Hélein [2] and Sandier–Serfaty [3]. Hence Theorem 2 can be viewed as their exact Finslerian extension.
@@ 5 Asymptotic Analysis and $\Gamma$–Convergence
Let us first recall that by the diamagnetic inequality (Lemma 5),
@@
Theorem 3 ($\Gamma$–convergence of $G_{F}$).
@@ 6 Numerical example
Discussion. The above analytic computations provide an explicit check that the Finslerian GL functional introduced in the paper recovers the expected anisotropic scaling in energy for simple topological configurations on the torus. The example is consistent with the general existence theory (Theorem 2) and with the $\Gamma$–-convergence statement (Theorem 3): energy concentration for sequences with growing winding or for configurations forcing zeros of $\psi$ would, after the standard $|\log\varepsilon|$–-rescaling, lead to limiting energies proportional to the Finsler length of the corresponding vortex currents; in this simple constant modulus family the energy is entirely carried by the phase gradient and computed above in closed form.
@@ 7 Concluding
In this work we have extended the classical Ginzburg–Landau theory to the setting of general Finsler manifolds. Starting from the analytic preliminaries in Section 2, we defined the anisotropic functional $G_{F}[\psi,A]$ in (16), proved its gauge invariance and well-posedness, and established the existence of minimizers (Theorem 2) by direct variational methods. Finally, in Section 5, we derived the full $\Gamma$–limit of the Finslerian energies as $\varepsilon\to 0$, showing that the limiting functional is the Finsler length of the vortex current.
@@
  • The Finsler diamagnetic inequality (Lemma 5) provides a geometric generalization of Kato’s inequality, valid for arbitrary convex co-metrics.
  • The variational proof of existence (Theorem 2) uses only the intrinsic Finsler Sobolev structure, avoiding Euclidean embeddings or local coordinates.
  • The $\Gamma$–limit (Theorem 3) identifies the limiting vortex energy with the anisotropic Finsler length $\int_{\Sigma}F(x,\nu_{\Sigma})\,d\mathcal{H}^{n-2}$, extending the isotropic theory of Bethuel–Brezis–Hélein and Sandier–Serfaty to arbitrary anisotropies.